# InferLog: Accelerating LLM Inference for Online Log Parsing via ICL-oriented Prefix Caching

### v1

@@ 6. Discussion
Recent LLM-based log parsers(Xu et al., 2024b; Jiang et al., 2024a; Huang et al., 2025; Wu et al., 2025; Xiao et al., 2024) design log parsing algorithm such as in-context learning, template caching or clustering, aiming to improve parsing accuracy. Conversely, InferLog targets the underlying LLM inference system, it acts as an intermediate layer between log parsers and the LLM inference system, InferLog is agnostic to the upstream parsers and seamlessly integrates with them to improve inference efficiency via prefix KV cache reusing and configuration tuning without compromising accuracy.
InferLog efficiently accelerates log parsing task than other LLM inference workloads.
Although methods like KV cache reusing and configuration tuning in InferLog can accelerate LLM inference in other contexts, there are two main features in log parsing make InferLog specifically suitable for log parsing and may lead to suboptimal performance in other scenarios.
Firstly, the order of in-context learning examples in log parsing has little impact on InferLog’s results. This is mainly because log parsing exhibits greater predictability and stability in its output structure, focusing on extracting log templates, so the model’s output is less affected by the order of examples. In contrast, in long document question answering, the relative position of information within the document is more critical for accurate answers(Jin et al., 2024b; Liu et al., 2023). Secondly, configuration tuning process requires a stable mapping between configuration and metrics. Log parsing task outputs stable log templates with predictable token length, creating a foundation for configuration tuning. However, in general workloads, LLM output length is dynamic, disrupting tuning model convergence.
@@ 6.1. Threats to Validity
Subject Systems. We conduct experiments on the vLLM inference system, due to its high popularity. We also compare the inference performance across various open-source LLMs. In the future, we plan to evaluate our method on more inference systems and models. InferLog only requires the target inference system and LLMs to support prefix caching, without the need for system modifications. Overall, InferLog can efficiently perform log parsing tasks on locally deployed open-source inference systems and LLMs without privacy concerns.
Randomness. Randomness may affect the performance through two aspects:(i) the randomness of LLM output and (ii) the randomness introduced during the performance testing of the system, particularly under online concurrent requests where task scheduling introduces variability in system performance. To mitigate the former threat, we set the temperature to 0, guaranteeing consistent outputs for identical input text. To mitigate the latter threat, each experiment was repeated three times for every experimental setting, and we report the the average results as the final outcome.
@@ 7. Related Work
LLM-based Log Parsing. Log parsing is a critical preliminary step for various log analysis tasks. Therefore, numerous efforts have been made to achieve accurate log parsing(He et al., 2017; Du and Li, 2016; Jiang et al., 2008; Messaoudi et al., 2018; Yu et al., 2023; Shima, 2016; Tang et al., 2011). With the rise LLMs, a series of LLM-based log parsers have been developed to achieve more effective log parsing(Xu et al., 2024b; Jiang et al., 2024a; Huang et al., 2025; Wu et al., 2025; Ma et al., 2024a, b; Zhong et al., 2024; Xiao et al., 2024). These LLM-based log parsers leverage fine-tuning(Ma et al., 2024a), in-context learning(Jiang et al., 2024a; Xu et al., 2024b; Wu et al., 2025)or log clustering analysis(Xiao et al., 2024; Wu et al., 2025) to specialize LLMs for log parsing tasks. Considering high invocation cost, some efforts have been made to achieve cost-effective LLM-based log parser. LILAC(Jiang et al., 2024a) enhances the efficiency of LLM-based log parsing by incorporating an adaptive parsing cache that stores parsing results, Logbatcher(Xiao et al., 2024) and LUNAR(Huang et al., 2025) partition logs and only queries LLM once for logs belonging to the same cluster. InferLog identifies the performance bottleneck in LLM inference under concurrent requests and optimized by ICL-oriented prefix caching, orthogonal to existing techniques for reducing query times.
@@ 8. Conclusion
In this work, we propose InferLog, a novel framework for optimizing LLM inference in online log parsing tasks. Based on the characteristics of LLM-based log parsing tasks and LLM inference performance, InferLog employs PAIR that consisted of prefix-based matching, modifying and reordering to refine ICL examples, thereby enhancing prefix KV cache hit rates. InferLog further combines attention-augmented meta-learning with SMBO for rapid, tailored configuration optimization. Experiments confirm its effectiveness in enhancing inference efficiency while preserving log parsing quality. The source code and data is available at https://github.com/wiluen/InferLog.

### v3

@@ 6.1. Practicality and Specialization
Recent LLM-based log parsers(Xu et al., 2024b; Jiang et al., 2024a; Huang et al., 2025; Wu et al., 2025; Xiao et al., 2024; Zhong et al., 2024) design log parsing algorithms such as in-context learning, template caching or clustering, aiming to improve parsing accuracy. Conversely, InferLog targets the underlying LLM inference system, it acts as an intermediate layer between log parsers and the LLM inference system, InferLog is agnostic to log parsers and seamlessly integrates with them to improve inference efficiency via prefix KV cache reusing and configuration tuning without compromising accuracy.
InferLog achieves higher efficiency in log parsing than in other LLM inference workloads.
Although methods like KV cache reusing and configuration tuning in InferLog can accelerate LLM inference in other contexts, there are two main features in log parsing make InferLog specifically suitable for log parsing and may lead to performance degradation in other scenarios.
Firstly, the order of ICL examples in log parsing has little impact on InferLog’s results. This is mainly because log parsing exhibits greater predictability and stability in its output structure, focusing on extracting log templates, so the model’s output is less affected by the order of examples. In contrast, in RAG or document-based inference, the relative index of information within the document is more critical for accurate answers(Jin et al., 2024b; Liu et al., 2023). Secondly, configuration tuning process requires a stable mapping between configuration and performance metrics. Log parsing task outputs stable log templates with predictable token length, creating a foundation for configuration tuning. However, in general workloads, the output lengths of LLMs are dynamic(e.g., varying with input context or task requirements), disrupting tuning model convergence.
@@ 6.2. Threats to Validity
Subject Systems. We conduct experiments on the vLLM inference system, due to its high popularity. We also compare the inference performance across various open-source LLMs. InferLog only requires the target inference system and LLMs to support prefix caching, without the need for system modifications. Overall, InferLog can efficiently perform log parsing tasks on locally deployed systems and open-source LLMs without privacy concerns.
Randomness. Randomness may affect the performance through two aspects: the randomness of LLMs output and the randomness introduced during the performance testing of the system, particularly under online concurrent requests where task scheduling introduces variability in system performance. To mitigate the former threat, we set the temperature to 0, guaranteeing consistent outputs for identical input text. To mitigate the latter threat, each experiment was repeated three times for every experimental setting, and we report the the average results as the final outcome.
@@ 7. Related Work
LLM-based Log Parsing. Log parsing is a critical preliminary step for various log analysis tasks. Therefore, numerous efforts have been made to achieve accurate log parsing(He et al., 2017; Du and Li, 2016; Jiang et al., 2008; Messaoudi et al., 2018; Yu et al., 2023; Shima, 2016; Tang et al., 2011). With the rise LLMs, a series of LLM-based log parsers have been developed to achieve more effective log parsing(Xu et al., 2024b; Jiang et al., 2024a; Huang et al., 2025; Wu et al., 2025; Ma et al., 2024a, b; Zhong et al., 2024; Xiao et al., 2024). These LLM-based log parsers leverage fine-tuning(Ma et al., 2024a), in-context learning(Jiang et al., 2024a; Xu et al., 2024b; Wu et al., 2025)or log clustering(Xiao et al., 2024; Wu et al., 2025) to specialize LLMs for log parsing tasks. Considering high invocation costs, some efforts have been made to achieve cost-effective LLM-based log parsers. LILAC(Jiang et al., 2024a) enhances the efficiency of LLM-based log parsing by incorporating an adaptive parsing cache that stores log templates, Logbatcher(Xiao et al., 2024) and LUNAR(Huang et al., 2025) partition logs and only queries LLM once for logs belonging to the same cluster. InferLog identifies the performance bottleneck in LLM inference under concurrent requests and optimized by ICL-oriented prefix caching, orthogonal to existing techniques for reducing the number of queries.
@@ 8. Conclusion
In this work, we propose InferLog, a novel framework of optimizing LLM inference for online log parsing of software systems. Based on the characteristics of LLM-based log parsing tasks and LLM inference performance, InferLog employs PAIR that consisted of prefix-based matching, modifying and reordering to refine ICL examples, thereby enhancing prefix KV cache hit rates. InferLog further combines attention-augmented meta-learning with SMBO for rapid, tailored configuration optimization. Experiments confirm its effectiveness in enhancing inference efficiency while preserving log parsing quality.
Artifact Availability. The source code and data is available at https://github.com/wiluen/InferLog.
9. Acknowledgments
The authors would like to thank the reviewers for their valuable comments and suggestions. This work was supported in part by National Key Research and Development Program of China (Grant Number: 2024YFB4505904), the National Natural Science Foundation of China under Grant 62272495 and the Guangdong Basic and Applied Basic Research Foundation under Grant 2023B1515020054. The corresponding author is Pengfei Chen.